Student X titrated 250cm3 of 0.1M with 0.1M HCl, using methyl orange as indicator. Student Y carried out the same exercise but used phenolphthalein as indicator which of the following statements about the titration is true?
a.the titre values obtained from the two titrations
b.the titre value of the titration by X is twice that by Y
c.the titre value of the titration by X is half that by Y
d.CO2 was released in the reactions in both titrations
1
Expert's answer
2020-06-29T07:44:45-0400

The reactions are

Na2CO3 + HCl = NaHCO3 + HCl

NaHCO3 + HCl = NaCl + H2O + CO2

Phenophthalein will give point up-to completion of first reaction and methyl orange will give end point after both reactions.

So if student rotate 250 Cm3 of 0.1 M with HCl using Phenolphthalein then

Moles of Na2CO3 = moles of HCl

250*0.1 = 0.1* x

So x = 250 ml.

Now again with methyl orange

Moles of Na2CO3 + mole of NaHCO3 = moles of HCl*2

250*0.1 = 0.1*x*2

X= 125 ml.

Hence option b is correct.

MiliMoles = M*Vin ml.

Due to NaHCO3 medium is weakly basic so Phenolphthalein is suitable till here and due to H2CO3 medium is slightly acidic so methyl orange is suitable till here.

H2CO3 means H2O and CO2.